OpenStudy (anonymous):

Which trinomial cannot be factored using integers? A) x^2-2x-63 B) x^2+16x+63 C) x^2-3x-3 D) x^2+5x-104

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (danjs):

do you know the term under the root in the quadratic formula b^2-4*a*c

OpenStudy (danjs):

y = a*x^2 + b*x + c \[x = \frac{ -b \pm \sqrt{b^2-4*a*c} }{ 2*a }\] calculate the root b^2 - 4 * a * c term calculate that for each one, see if you spot anything

OpenStudy (danjs):

or calculate the whole thing and see which values of x are not integers (whole numbers including the negatives)